# Mode-conditioned music learning and composition: a spiking neural network inspired by neuroscience and psychology

### v1

@@ Abstract
Keywords Symbolic music  $\cdot$ Brain-inspired Spiking neural network  $\cdot$ Mode perception  $\cdot$ Music learning  $\cdot$ Music generation
@@ 1 Introduction
Music learning and generation have long fascinated researchers across various disciplines, including neuroscience, psychology, and artificial intelligence. In particular, the artificial intelligence field has witnessed significant advancements in recent years, driven by the integration of deep learning techniques, large-scale models, and novel neural network architectures[1, 2]. With the emergence of OpenAI’s Sora [3], the techniques of music learning and generation seem so powerful that they are capable of producing compositions that rival human-created music in complexity and emotional depth. However, despite these impressive advancements, the biggest challenge remains: current techniques lack the ability to understand and generate music on a cognitive level. Fig. 1 exhibits the Macro comparison of music learning and composition between human being and current deep learning techniques.
Figure 1: The comparison of music composition between human, traditional artificial intelligent algorithms, and the brain-inspired models.
Traditional AI
Since the first music piece was created by a computer at Bell Laboratories in 1957[4], the techniques of automatic music generation have developed over past six decades. Symbolic music, which represents music through notational symbols like pitch, duration, and instrumentation, captures detailed musical elements such as melody, harmony, and structure. These features make it particularly well-suited for deep music analysis. Recent years, deep learning has emerged as a powerful tool for symbolic music, with numerous landmark studies making significiant contributions to this field. In 1998, Todd[5, 6] firstly used recurrent neural network(RNN) to generate music melodies. Later in 1994, Mozer developed CONCERT[7, 1] system, also based on RNN, to generate melodies by conditioning on pitch, duration, and harmonic chord accompaniment. Eck and Schmidhuber were the first to employ Long-Short Term Memory(LSTM) to overcome the vanishing gradient problems, creating blues music with this method[8]. Since then, RNNs, particularly their LSTM variants, have become one of the most popular models for music learning and generation. Boulanger-Lewandowski et al.[9] made further advancements by using RNN-RBM model to learn harmonic and rhythmic rules and polyphonic music. Google Brain’s Magenta project later developed MelodyRNN[10] model, incorporating LookBackRNN and attention mechanisms to better capture the long-term structure of music. Additionally, DeepBach[11] was designed to generate four-part harmonizations in the style of Johann Sebastian Bach’s chorales, using a combination of Bi-direction LSTM to capture both the past and future context of music sequences. Besides, models based on Variational AutoEncoder(VAEs) or Generative Adversarial Networks (GANs) architecture have also been explored for analyzing and generating music. MusicVAE[12], for instance, uses hierarchical VAE structure to model both monophonic and polyphonic music sequences, generating relative consistent music pieces. Wang and et al.[13] proposed PianoTree VAE, which is designed as a new tree-structure extension model to generated polyphonic counterpoint music. Wei and Xia[14] proposed a EC2-VAE to learn long-term symbolic representations through contextual constrains. MidiNet[15] employs a generative adversarial network to generate more interesting multi-track music than MelodyRNN, which based on a novel conditional mechanism for generating melodies that follow a chord sequence, or by the previous bars. Dong et al.[16] introduced MuseGAN, which generates five-tracks music pieces within the GANs, including three models. Researchers also utilized GANs to generate different styles of music, such as Jazz[17]. Transformers-based models recently have gained popularity in the music generation domain. Huang et al.[18] explored the language model Transformer XL to generated popular piano music, while MuseNet[19] can produce several minutes music with various instruments and genres based on GPT-2. With the widespread use of large language models (LLMs), Zhao[20] proposes Adversarial-MidiBERT, to achieve four music understanding tasks. Xu et al.[21] developed a text-conditioned music generation model based on the LLM-enhanced MetaScore. In summary, as shown in the green part of Fig.1, current computational models predominantly rely on large datasets and parameter-heavy architectures. This reliance presents a significant limitation: it renders the learning process increasingly complex, diverging significantly from the intuitive and explainable learning and creative processes exhibited by human beings.
Human Being
Music is intrinsic to human nature, and researchers of cognitive neuroscience have shown that it is a so complex cognitive activity requiring multiple brain areas to collaborate [22]. As shown in the yellow part of Fig.1, the perception and creation of music involve intricate processes that engage various regions of the brain, including those responsible for auditory processing, memory, motor control, emotional regulation and etc. Studies have shown that the primary auditory cortex(A1) has a topographical structure that maps the cochlea. The organization implies that neural populations in different regions response to different frequencies of sound[23, 24, 25, 26]. Theses findings clearly demonstrated that the encoding manner of the sound in brain. However, through there are no clear and consistent findings regarding how the brain encodes and process temporal information in music perception, some researches have found that a significant number of cells in the medial premotor cortex (MPC) are sensitive to a range of signal durations, with their preferred durations spanning intervals from several hundred milliseconds[27, 28], and "Time Cells" [29] may encode moments between temporal events and locations. These findings may be helpful for music time perception. Besides, one significant area of exploration is the role of memory in music cognition. Memory allows for the recognition of familiar melodies, the anticipation of musical patterns, and the contextual understanding of musical pieces. This involves both short-term memory, which helps in processing and interpreting ongoing musical input, and long-term memory, which stores musical knowledge and experiences accumulated over time. Researchers have found that, hippocampus, medial temporal lobe(MTL) as well as prefrontal cortex(PFC) are deeply involved in temporal context memory establishment[30, 31, 32, 33]. Prior knowledge also exerts critical influence on memory and is one of the prominent factor guiding the learning behaviors across human lifespan. Studies have highlighted the significant roles of the medial prefrontal cortex (mPFC) and the hippocampus (HC) in the formation of prior knowledge and its utilization during the successful encoding, consolidation, and retrieval of memories[34].
Additionally, principles of music theory provide the powerful supports to ensure that the music learned and generated remains reasonable and consistent. Modal theory, especially, which deals with the use of different scales and modes in music composition, offers a framework for capturing the tonal and harmonic nuances that are essential to many musical traditions. In the domain of music psychology, one of the most important model is the Krumhansl-Schmuckler Key-Finding Model which was developed by Carol Krumhansl and Edward Schmuckler in the early 1990s[35, 36]. This model developed the idea of tonal hierarchies through empirical experiments, where listeners rated the importance or fit of each note within a given key. For each major and minor key, Krumhansl constructed a key profile, which represents the relative importance or frequency of the 12 pitch classes (C, C#, D, D#, etc.). Each profile consists of values based on the experimental data from listeners. The model compares the pitch-class distribution from the musical input to each of the 24 key profiles (12 major and 12 minor). It uses correlation to determine how well the pitch-class distribution matches each key profile. The key that yields the highest correlation is considered the most likely key of the piece. This model plays a significant role in understanding tonal perception in Western music and has influenced computational approaches in music theory and cognitive science.
Brain-inspired Models
Facing the gap between technical proficiency and true musical comprehension, brain-inspired methods integrate brain mechanisms with the strengths of traditional AI models to enhance the model’s cognitive abilities, as shown in the red part of Fig.1. Some researchers[37] have utilized fMRI responses of human brain activity to music stimuli as embeddings for deep learning models. However, the primary goal of such studies has been to reconstruct music, rather than to explore the deeper aspects of musical understanding or creative generation. Spiking Neural Networks (SNNs), whose neurons and learning principles more closely resemble those of the human brain, offer a potential solution to the current challenges. Q. L. and colleagues proposed a spiking neural network inspired by cortical structures to memorize music [38], and to generate stylistic [39] and emotional [40] musical melodies. However, these studies primarily focus on monophonic melody generation and lack the integration of music theory. Overall, the application of brain-inspired spiking neural networks for learning and generating music remains rare. Addressing the current challenges and the scarcity of research in this area, this paper proposes a brain-inspired SNN model that learns and generates multi-track music by combining brain mechanisms, music psychology findings, and Western mode and key theories. Fig.2 shows the total architecture of the model. The contributions of this paper are as follows:
@@
  • We build a multi-area collaborative model [^1] We released our code in https://github.com/BrainCog-X/Brain-Cog/tree/main/examples/Knowledge_Representation_and_Reasoning/musicMemory/task based on a Spiking Neural Network (SNN) on the open platform BrainCog [40], inspired by neuroscience findings, to perceive, learn, and create music. We employ the Izhikevich neuron model within the SNN and utilize synaptic creation along with the Spike-Timing-Dependent Plasticity (STDP) learning rule to facilitate dynamic circuit evolution, which enhances the network’s adaptability. This mechanism closely mirrors the ways in which learning and memory are implemented in the human brain, providing a biologically plausible approach to music learning and generation.
@@
  • The model incorporates Western mode theory, neuroscientific mechanisms, and insights on mode perception from music psychologists, areas that have been insufficiently explored by existing computational algorithms. By integrating mode theory, our model aims to capture tonal characteristics and exhibit a connection framework similar to the well-known Krumhansl-Schmuckler pitch profile model. This approach enhances the musicality and authenticity of the generated compositions. For understanding the model clearly, the musical terms used in this paper are listed in Table.1
@@
  • Our evaluation framework includes both model evaluation and generation quantitative assessments. The model evaluation focus on the comparation with the model framework and the psychology model. The generation quantitative metrics measure involve the musicality, and aesthetic quality of the generated music. The results shows that the proposed model is consistent with the Krumhansl-Schmuckler psychological model, and the qualitative assessment of the generated samples demonstrates that the generated pieces effectively capture the intended tonal characteristics of different modes and keys, further supporting the efficacy of our approach in bridging the gap between artificial intelligence and human-like musicality.
@@
2 Results
2.1 Dataset Description
J.S. Bach’s four-part chorales(Bach): To make the model ’listen’ more music works, we employ the J.S. Bach’s four-part chorales dataset, which is available through the Music21 python package [41]. It includes 408 famous chorales in musicxml format, with 219 in major keys and 189 in minor keys, excluding Db and Gb major, as well as Db, Eb, and Ab minor. A summary of the key distributions is provided in Table.2
Table 2: The number of keys in SHTE and Bach datasets
Keys C C#/Db D D#/Eb E F F#/Gb G G#/Ab A A#/Bb B
Major SHTE 6 4 11 8 10 13 1 8 8 11 13 3
Bach 22 - 32 9 10 28 - 59 1 32 25 1
Minor SHTE 9 6 10 - 11 7 17 17 - 9 2 9
Bach 10 - 27 - 20 2 7 44 - 49 2 28
2.2 Experiments
2.2.1 Our Model VS KS Model
    $PSC_{k}=\sum^{4}_{j=1}\sum^{128}_{r\%12=k}\sum^{N}_{s}c_{k,jrs}$ (1)
For the minor mode(Fig.3(B)(D)), the PSCs and PASWs also play the most important role at tonic(C). However, in the KS model, the scores for mediant(Eb)$KS_{m}inor(Eb)=0.75$ are higher than those for the dominant(G)$KS_{m}inor(G)=0.58$. In contrast, the PSC and PASW values in the SHTE dataset($PSC_{SHTE}(Eb)=0.52$, $PASW_{SHTE}(Eb)=0.45$) and Bach dataset ($PSC_{Bach}(Eb)=0.74$, $PASW_{Bach}(Eb)=0.58$) are lower than those at the dominant(G). This discrepancy can be attributed to the key and necessary role of the dominant chord in the cadences and its frequent use in both datasets. Besides, compared to KS model in SHTE dataset(3(B)), these two metrics for the leading tone(B) show an ascending trend, This is due to the teaching characteristics of the SHTE dataset, which includes more exercises focused on learning chords.
2.2.2 Quality Evaluation And Analysis Of Generation
To further assess and analyze the quality of the generated music pieces, we employ a comprehensive objective method which is inspired by the method proposed by Yang and Lerch[42]. Several features are extracted from the generated samples and two datasets involved in our model. The features include: Pitch Count(PC), the number of different pitches in a music piece; Pitch Class Histogram(PCH), the histogram of 12 pitch classes; Diatonic Pitch Rate(DPR), the proportion of pitches in a music piece that belongs to the key’s diatonic scale; Pitch Range(PR), the pitch range between the highest and the lowest pitch in semitones within a sample; Pitch Interval(PI), average interval between two pitch neighbors, PCTM, the pitch class transition matrix.
We randomly select 50 samples of our training datasets SHTE and the J.S. Bach’s(JSB), respectively. Similarly, we randomly select 50 samples of our generated 8-bars long music pieces as the test dataset, covering different keys for testing. We compute the features mentioned above for these three datasets. Both absolute and relative measurements are discussed for these features mentioned above. The absolute measurements specially refer to the mean and the standard deviation, focusing on the characteristics within each dataset. On the ohter hand, the relative measurements emphasize the features across datasets, with the Kullback–Leibler divergence(KLD) and overlapped area(OA) being employed to evaluate the similarities between the generated music pieces and the two datasets. The experimental results are summarized in Table. 3.
SHTE Bach Generated Data
Mean STD Mean STD Mean STD KLD OA KLD OA
PC 8.00 1.79 8.40 2.06 7.70 1.27 0.13 0.67 0.11 0.71
PCH - - - - - - 0.16 0.59 0.71 0.45
DPR 0.96 0.04 0.93 0.04 0.86 0.04 0.02 0.78 0.03 0.83
PR 11.9 3.51 11.7 2.23 9.4 2.2 0.05 0.74 0.01 0.82
PI 2.51 1.01 1.92 0.41 2.33 0.32 0.11 0.55 0.04 0.71
PCTM - - - - - - 0.09 0.45 0.19 0.28
Referring to the evaluation results, the analysis are as follows:
  • Besides, for the PCH and PCTM features,The KLD of PCH is lower for SHTE ($0.16$) than for Bach ($0.71$), meaning the pitch class histogram of the generated data is more aligned with SHTE. The OA values for PCH are $0.59$ (SHTE) and $0.45$ (Bach), also suggest that the generated data resembles SHTE more in pitch class distribution patterns. Similarly, KLD values for PCTM are higher with Bach ($0.19$) than with SHTE ($0.09$), and OA values are $0.45$ (SHTE) and $0.28$ (Bach), suggesting that the pitch transitions of generated data more closely resemble SHTE.
In summary, the generated data aligns more closely with Bach’s melodic characteristics while adopting some unique features from SHTE, such as a tighter pitch class spread and more distinct transition patterns, contributing to a balance of traditional structure and innovative variation. The results further indicate that the proposed model effectively captures both the tonality characteristics and the melody adaptability needed for generating diverse and musical content. The model’s ability to blend well-established tonal principles with novel musical expressions underscores its potential for generating musically meaningful compositions across various styles and datasets. This capability is essential for our future work on harmonic learning, emotion perception, and stylistic composition.
@@ 3.1.2 Sequential Memory Subsystem
As illustrated in Figure 2, the pitch and duration subnetworks share the same internal connection structure. The synaptic connections between adjacent layers and across layers are excitatory and fully connected. Meanwhile, the introduction of synaptic plasticity and transmission delay plays a crucial role in the learning process. Synaptic plasticity allows the connection weights between neurons to be adjusted based on experience and learning, mirroring the adaptive capabilities of biological neural networks. Transmission delay, which accounts for the time it takes for signals to travel between neurons, is also incorporated to ensure that the timing of neural firing patterns is accurately represented. Connections within the same layer are inhibitory. For further details about the sequential memory system, see previous works [38, 39].

### v2

@@ Abstract
keywords:
[…]
@@ 1 Introduction
Music learning and generation have long fascinated researchers across various disciplines, including neuroscience, psychology, and artificial intelligence. In particular, the artificial intelligence field has witnessed significant advancements in recent years, driven by the integration of deep learning techniques, large-scale models, and novel neural network architectures Briot et al. (2020); Hernandez-Olivan et al. (2022). With the emergence of OpenAI’s Sora OpenAI (2024), the techniques of music learning and generation seem so powerful that they are capable of producing compositions that rival human-created music in complexity and emotional depth. However, despite these impressive advancements, the essential challenge also remains: current techniques lack the ability to understand and generate music on a cognitive level. This challenge gives rise to two key problems: 1) current computational model do not have the real ability of music perception and understanding, meaning they cannot interpret or comprehend music in the way humans do, 2) results of music learning and generating consequently fail to capture the depth, nuance, and creativity process inherent to human composers, finally limiting the consistency and musicality of generated results.
Music is intrinsic to human nature, and researchers of cognitive neuroscience have shown that it is a so complex cognitive activity requiring multiple brain areas to collaborate Koelsch (2012). The perception and creation of music involve intricate processes that engage various regions of the brain. The primary auditory cortex (A1) plays a crucial role in sound processing, encoding the real sound signals to neural electrophysiological signals Hodges and Thaut (2019). The memory system is responsible for storing the sequential information of ordered notes, which is fundamental for recognizing familiar melodies, recalling harmonic progressions, and understanding complex musical forms. Prefrontal cortex, which is the most important region to represent concept and knowledge of the world, plays a crucial role in guiding learning and creating behaviors of human Bi (2021). Additionally, music has a profound emotional impact, with the emotion pathway, including the amygdala and orbitofrontal cortex playing central roles in encoding and generating emotional responses to music Hodges and Thaut (2019). The integration of these systems enables humans to experience and create music in a deeply meaningful and expressive way. Understanding these neural mechanisms provides a foundation for developing models that aim to make AI models more cognitive inspired, further improving their ability to generate music with enhanced emotional depth, coherence, and human-like creativity.
Mode theory and key play a central role in music composition, serving as the foundation for establishing tonal centers and guiding the melodic and harmonic structure of a piece. Together, they provide a coherent framework that shapes the emotional character and direction of music. The interplay between mode and key is crucial in creating process, influencing how listeners perceive mood and movement in a composition. Despite its importance, this fundamental theory is often overlooked by current AI music generation models. Most existing models adopt a simplified approach by aligning all keys to either the C major or minor modes, disregarding the subtle tonal colors and unique relationships between notes in different modes. However, researchers in music psychology domain have found that humans possess innate key perception mechanisms, enabling them to distinguish and interpret different keys and modes based on tonal hierarchies. To explain this cognitive process, one of the most well-known models, the Krumhansl-Schmuckler Model(KS), have been proposed. This model demonstrates how listeners perceive tonal centers and assign hierarchical importance to pitch classes within a key, providing a quantitative framework for understanding human key perception. Despite these advances in cognitive music theory, current AI models rarely incorporate such insights, limiting their ability to generate music that reflects the nuanced key perception and tonal complexity inherent in human compositions. Consequently, there remains a significant gap between AI-generated music and the complexity of human creativity.
Inspired by the mechanisms of human music cognition, and addressing the limitations of current techniques in capturing the cognitive depth of music perception and composition, this paper proposes a multi-region collaborated brain-inspired spiking neural network that learns and generates multi-track music by integrating key principles from brain mechanisms, music psychology findings, and Western mode and key theories. The contributions of this paper are as follows:
@@
  • We build a multi-area collaborative model [^1] We released our code in https://github.com/BrainCog-X/Brain-Cog/tree/main/examples/Knowledge_Representation_and_Reasoning/musicMemory/task based on a Spiking Neural Network (SNN) on the open platform BrainCog Zeng et al. (2023), inspired by neuroscience findings, to perceive, learn, and create four-part music based on the symbolic music form. We employ the Izhikevich neuron model within the SNN and utilize synaptic creation along with the Spike-Timing-Dependent Plasticity (STDP) learning rule to facilitate dynamic circuit evolution, which enhances the network’s adaptability.
@@
  • The model incorporates Western mode theory, inspired by the mechanisms of representation of prior knowledge of prefrontal cortex, building a hierarchical subsystem to represent and learn modes, keys and the relationship between them. By integrating mode theory, our model aims to capture tonal characteristics and exhibit a connection framework similar to the well-known Krumhansl-Schmuckler pitch profile model. For understanding the model clearly, the musical terms used in this paper are listed in Table.1
@@
  • We build an evaluation framework including both model evaluation and generation quantitative assessments. The results show that the architecture of our proposed model is consistent with the Krumhansl-Schmuckler psychological model, and the qualitative assessment of the generated samples demonstrates that the generated pieces effectively capture the intended tonal characteristics of different modes and keys, further supporting the efficacy of our approach in bridging the gap between artificial intelligence and human-like musicality.
2 Related Works
2.1 Traditional AI
Symbolic music learning primarily focuses on music analysis. Current popular AI models on for symbolic music learning include Recurrent Neural Networks (RNNs), Variational AutoEncoders (VAEs), the Generative Adversarial Networks (GANs), Transformers and etc. Previous Works using RNNs and LSTMs targeted melody generation Todd (1989); Ji et al. (2023); Mozer (1994); Briot et al. (2020), harmonic Eck and Schmidhuber (2002); Waite (2016); Boulanger-Lewandowski et al. (2012), chord accompaniment Eck and Schmidhuber (2002), polyphonic music Boulanger-Lewandowski et al. (2012) and stylistic variation Eck and Schmidhuber (2002); Hadjeres et al. (2017). VAEs and GANs have also been employed to improve polyphony Wang et al. (2020), style adaptation Trieu and Keller (2018), and multi-track generation Roberts et al. (2018); Shiqi Wei (2022); Yang et al. (2017); Dong et al. (2018), Recently, Transformers-based models recently have gained popularity in the music generation domain Huang and Yang (2020). Then large language models (LLMs) based works  Payne (2019); Zhao (2024); Xu et al. (2024) have further enhanced symbolic music generation. However, these computational models predominantly rely on large datasets and parameter-heavy architectures, overlooking the central role of mode and key theories in music composition. Furthermore, they diverge significantly from the intuitive, explainable, and cognitively driven learning and creative processes exhibited by human composers, resulting in outputs that lack the depth and coherence of human-generated music.
2.2 Neuroscience and Psychology
Studies have shown that the primary auditory cortex(A1) has a topographical structure that maps the cochlea. The organization implies that neural populations in different regions response to different frequencies of sound Vuust et al. (2022); Purves et al. (2001); Sankaran et al. (2024); McDermott and Oxenham (2008). Theses findings clearly demonstrated that the encoding manner of the sound in brain. However, through there are no clear and consistent findings regarding how the brain encodes and process temporal information in music perception, some researches have found that a significant number of cells in the medial premotor cortex (MPC) are sensitive to a range of signal durations, with their preferred durations spanning intervals from several hundred milliseconds Merchant et al. (2013); Gupta (2014), and "Time Cells"  MacDonald et al. (2011) may encode moments between temporal events and locations. These findings may be helpful for music time perception. Besides, one significant area of exploration is the role of memory in music cognition. Memory allows for the recognition of familiar melodies, the anticipation of musical patterns, and the contextual understanding of musical pieces. This involves both short-term memory, which helps in processing and interpreting ongoing musical input, and long-term memory, which stores musical knowledge and experiences accumulated over time. Researchers have found that, hippocampus, medial temporal lobe(MTL) as well as prefrontal cortex(PFC) are deeply involved in temporal context memory establishment Tubridy and Davachi (2011); McAndrews and Milner (1991); Meier et al. (2013); Jenkins and Ranganath (2010). Prior knowledge also exerts critical influence on memory and is one of the prominent factor guiding the learning behaviors across human lifespan. Studies have highlighted the significant roles of the medial prefrontal cortex (mPFC) and the hippocampus (HC) in the formation of prior knowledge and its utilization during the successful encoding, consolidation, and retrieval of memories Brod et al. (2013).
Additionally, the Krumhansl-Schmuckler Model Krumhansl (1990); Temperley (1999), which is one of the most important model in music psychology domain, developed the idea of tonal hierarchies through empirical experiments, where listeners rated the importance or fit of each note within a given key. For each major and minor key, Krumhansl constructed a key profile, which represents the relative importance or frequency of the 12 pitch classes (C, C#, D, D#, etc.). Each profile consists of values based on the experimental data from listeners. The model compares the pitch-class distribution from the musical input to each of the 24 key profiles (12 major and 12 minor). It uses correlation to determine how well the pitch-class distribution matches each key profile. The key that yields the highest correlation is considered the most likely key of the piece. This model plays a significant role in understanding tonal perception in Western music and has influenced computational approaches in music theory and cognitive science.
2.3 Brain-inspired Models
Facing the gap between technical proficiency and true musical comprehension, brain-inspired methods integrate brain mechanisms with the strengths of traditional AI models to enhance the model’s cognitive abilities. Some researchers Denk et al. (2023) have utilized fMRI responses of human brain activity to music stimuli as embeddings for deep learning models. However, the primary goal of such studies has been to reconstruct music, rather than to explore the deeper aspects of musical understanding or creative generation. Spiking Neural Networks (SNNs), whose neurons and learning principles more closely resemble those of the human brain, offer a potential solution to the current challenges. Q. L. and colleagues proposed a spiking neural network inspired by cortical structures to memorize music  Liang et al. (2020), and to generate stylistic  Liang and Zeng (2021) and emotional  Zeng et al. (2023) musical melodies. However, these studies primarily focus on monophonic melody generation and lack the integration of music theory. Overall, the application of brain-inspired spiking neural networks for learning and generating music remains rare.
@@ 3.1.2 Sequential Memory Subsystem
As illustrated in Figure 1, the pitch and duration subnetworks share the same internal connection structure. The synaptic connections between adjacent layers and across layers are excitatory and fully connected. Meanwhile, the introduction of synaptic plasticity and transmission delay plays a crucial role in the learning process. Synaptic plasticity allows the connection weights between neurons to be adjusted based on experience and learning, mirroring the adaptive capabilities of biological neural networks. Transmission delay, which accounts for the time it takes for signals to travel between neurons, is also incorporated to ensure that the timing of neural firing patterns is accurately represented. Connections within the same layer are inhibitory. For further details about the sequential memory system, see previous works Liang et al. (2020); Liang and Zeng (2021).